# MEGA: Memory-Efficient 4D Gaussian Splatting for Dynamic Scenes

### v1

@@ 3.2 Memory-efficient 4D Gaussian Splatting for Dynamic Scenes
Entropy-constrained Gaussian Deformation. For a specific time $t$, 4DGS (Yang et al., 2024a) presupposes that the sliced 4D Gaussians exhibit linear movement while their rotation and scale remain constant. This strict assumption simplifies the movement representation and forces the model to combine multiple extra Gaussians to present any complex non-linear motions. Moreover, the sliced 4D Gaussian introduces the temporal decay opacity $\sigma_{t}$. From its definition, it is analyzed that a Gaussian gradually appears as time $t$ approaches its temporal position $\mu_{t}$, peaks in opacity at $t=\mu_{t}$, and gradually diminishes in density as $t$ moves away from $\mu_{t}$. As shown in Fig. 4 (a), this limited temporal operation range results in more than 90% of Gaussians being excluded at each time, causing the model to densify a large amount of Gaussians for rendering high-quality scene.
@@
To address these limitations, we advocate for improving flexibility in the motion representation and geometric structure of each 4D Gaussian. Specifically, we introduce a temporal-viewpoint aware deformation predictor to enlarge the action range of Gaussians. The 4D Gaussian center $\bm{\mu}_{4D}$, view direction $\bm{d}_{i}$, and time $t$ are mapped to a high-dimensional space using a regular frequency positional encoding function $\gamma$ (Mildenhall et al., 2021), and then processed through a lightweight MLP network $\mathcal{F}_{\bm{\theta}}$ to predict the position deformation $\bm{m}_{\bm{\mu}_{4D}}^{t,v}\in\mathbb{R}^{4}$, scale deformation $\bm{m}_{\bm{s}_{4D}}^{t,v}\in\mathbb{R}^{4}$, and rotation deformations $\bm{m}_{\bm{q}_{l}}^{t,v}\in\mathbb{R}^{4},\bm{m}_{\bm{q}_{r}}^{t,v}\in\mathbb{R}^{4}$ as:
$(\bm{m}_{\bm{\mu}_{4D}}^{t,v},\bm{m}_{\bm{s}_{4D}}^{t,v},\bm{m}_{\bm{q}_{l}}^{t,v},\bm{m}_{\bm{q}_{r}}^{t,v})=\mathcal{F}_{\bm{\theta}}(\gamma({\rm sg}(\bm{\mu}_{4D})),\gamma({\rm sg}(\bm{d}_{v})),\gamma(t)),$ (4)
@@
$\bm{\mu}_{4D}^{t,v}=\bm{\mu}_{4D}\times\bm{m}_{\bm{\mu}_{4D}}^{t,v},\quad\bm{s}_{4D}^{t,v}=\bm{s}_{4D}\times\bm{m}_{\bm{s}_{4D}}^{t,v},\quad\bm{q}_{l}^{t,v}=\bm{q}_{l}\times\bm{m}_{\bm{q}_{l}}^{t,v},\quad\bm{q}_{r}^{t,v}=\bm{q}_{r}\times\bm{m}_{\bm{q}_{r}}^{t,v}.$ (5)
@@ 3.3 Training and Compression Pipeline
Loss Function. Following the original 4DGS (Yang et al., 2024a), we adopt the photometric loss $\mathcal{L}_{photo}$, consisting of $\mathcal{L}_{1}$ loss and structural similarity loss $\mathcal{L}_{ssim}$, to measure the distortion between the rendered image and ground truth image. By adding the loss for opacity regularization $\mathcal{L}_{opa}$, the overall loss $\mathcal{L}$ is defined as:
@@
Table 1: Quantitative comparison with various competitive baselines on the Technicolor Dataset. “Storage” refers to the total model size for 50 frames.
[…]
DyNeRF 31.80 - 0.0210 0.1400 0.02 30.00MB
HyperReel 32.70 0.0470 - 0.1090 4.00 60.00MB
Deformable 3DGS 30.95 0.0696 0.0353 0.1553 76.09 61.36MB
STG 33.35 0.0404 0.0187 0.0846 141.73 51.35MB
E-D3DGS 32.89 0.0494 0.0231 0.1114 79.14 56.07MB
4DGS 32.07 0.0535 0.0263 0.1189 55.26 6107.07MB
Ours 33.57 0.0442 0.0204 0.1014 83.14 32.45MB
@@ 4.1 Experimental Setup
Datasets. We evaluate the effectiveness of our method using two real-world benchmarks that are representative of various challenges in dynamic scene rendering: (1) Technicolor Light Field Dataset (Sabater et al., 2017): This dataset consists of multi-view video data captured by a time-synchronized 4$\times$4 camera rig. Following HyperReel (Attal et al., 2023), we exclude the camera at the second row, second column and evaluate on five scenes (Birthday, Fabien, Painter, Theater, and Trains) at 2048$\times$1088 full resolution. (2) Neural 3D Video Dataset (Neu3DV) (Li et al., 2022b): This dataset includes six indoor multi-view video scenes captured by 18 to 21 cameras, each at a resolution of 2704×2028 pixels. The scenes (Coffee Martini, Cook Spinach, Cut Roasted Beef, Flame Salmon, Flame Steak, Sear Steak) vary in duration and feature dynamic movements, some with multiple objects in motion. Consistent with existing practices (Li et al., 2022b; Yang et al., 2024a), evaluations are conducted at half resolution of 300-frame scenes.
Evaluation Metrics. To assess the quality of rendered videos, we utilize three popular image quality assessment metrics: Peak Signal-to-Noise Ratio (PSNR), Dissimilarity Structural Similarity Index Measure (DSSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018). PSNR quantifies the pixel color error between the rendered and original frames. DSSIM evaluates the perceived dissimilarity of the rendered image, while LPIPS measures the higher-level perceptual similarity using an AlexNet backbone (Krizhevsky et al., 2012). Given the inconsistency in DSSIM implementation noted across different methods (Fridovich-Keil et al., 2023; Attal et al., 2023), we follow Li et al. (2024) to distinguish DSSIM results into two categories: DSSIM1 and DSSIM2. DSSIM1 is calculated with a data range set to 1.0, based on the structural similarity function from the scikit-image library, whereas DSSIM uses a data range of 2.0. For rendering speed, we measure the performance in frames per second (FPS).
Table 2: Quantitative comparisons with various competitive baselines on the Neural 3D Video Dataset. “Storage” refers to the total model size for 300 frames. 1: Only report the result on the Flame Salmon scene. 2: Exclude the Coffee Martini scene. 3: These methods train each model with a 50-frame video sequence to prevent memory overflow, requiring six models to complete the overall evaluation.
@@
Neural Volume1 22.80 - 0.0620 0.2950 - -
LLFF1 23.24 - 0.0200 0.2350 - -
DyNeRF1 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 31.71 - 0.0140 0.0750 0.56 200.00MB
StreamRF 28.26 - - - 10.90 5310.00MB
NeRFPlayer3 30.69 0.0340 - 0.1110 0.05 5130.00MB
HyperReel 31.10 0.0360 - 0.0960 2.00 360.00MB
K-Planes 31.63 - 0.0180 - 0.30 311.00MB
MixVoxels-L 31.34 - 0.0170 0.0960 37.70 500.00MB
MixVoxels-X 31.73 - 0.0150 0.0640 4.60 500.00MB
Dynamic 3DGS 30.46 0.0350 0.0190 0.0990 460.00 2772.00MB
C-D3DGS 30.46 - - 0.1500 118.00 338.00MB
Deformable 3DGS 30.98 0.0331 0.0191 0.0594 29.62 32.64MB
E-D3DGS 31.20 0.0259 0.0151 0.0304 69.70 40.20MB
STG3 32.04 0.0261 0.0145 0.0440 273.47 175.35MB
4DGS 31.57 0.0290 0.0164 0.0573 96.69 3128.00MB
@@
Baselines. As we introduce MEGA, a novel method for compressing 4DGS (Yang et al., 2024a), our primary comparison focuses on the baseline 4DGS method. Additionally, we benchmark MEGA against a range of NeRF-based baselines, including DyNeRF (Li et al., 2022b), HyperReel (Attal et al., 2023), Neural Volume (Lombardi et al., 2019), LLFF (Mildenhall et al., 2019), HexPlane (Cao & Johnson, 2023), NeRFPlayer (Song et al., 2023), MixVoxels (Wang et al., 2023), and K-Planes (Fridovich-Keil et al., 2023). Other recent competitive Gaussian-based methods are also considered in our comparisons, including Dynamic 3DGS (Luiten et al., 2024), C-D3DGS (Katsumata et al., 2024), Deformable 3DGS (Wu et al., 2024), E-D3DGS (Bae et al., 2024), and STG (Li et al., 2024). The numerical results of Deformable 3DGS, E-D3DGS, STG, and 4DGS are produced by running their released codes on a single NVIDIA A800 GPU, while results for other baselines are from their original papers.
@@
Implementation Details. We train our MEGA model over 30k iterations and stop densification at the midpoint. We use the Adam optimizer with a batch size of one, adopting the hyperparameter settings from the original 4DGS (Yang et al., 2024a) framework, including loss weight, learning rate, and threshold parameters. When rendering the view at time $t$, we filter out those Gaussians with $\sigma(t)\leq 0.05$. To ensure stable training of our deformation predictor, we introduce weight regularization and set it at $1e^{-6}$. The learning rate of the deformation predictor undergoes exponential decay, starting from $8e^{-4}$ and reducing to $1.6e^{-6}$. For the AC color predictor, we start with an initial learning rate of 0.01, incorporating a 100-step warm-up phase. Subsequently, its learning rate is decreased by a factor of three at the 5k, 15k, and 25k steps. Regarding the hyper-parameters in the loss function, we set $\lambda$ and $\kappa$ as 0.2 and 0.0005, respectively, to balance the contributions of different components.
4.2 Experimental Results
Table 1 details a quantitative evaluation of our MEGA method on the Technicolor dataset. Notably, our method surpasses the main baseline 4DGS (Yang et al., 2024a), with PSNR, DSSIM1, DSSIM2, and LPIPS improvements by 1.2dB, 0.01, 0.006, and 0.018, respectively. Meanwhile, it significantly reduces storage requirements, achieving a 190$\times$ compactness and improving rendering speed by 50%. When compared with the NeRF-based method HyperReel (Attal et al., 2023), MEGA achieves a substantial improvement in representation, with an increase of about 0.87dB in PSNR and a 20$\times$ faster rendering speed, while halving the storage overhead. Moreover, our MEGA records a 0.22dB gain in visual fidelity over the state-of-the-art (SOTA) Gaussian-based method STG (Li et al., 2024), and reduces storage overhead by 40%. Fig. 5 offers qualitative comparisons for the Theater and Painter scenes, demonstrating that our results contain more vivid details and provide artifact-less rendering. More visual comparisons are available in Appendix A.
Besides, we report the quantitative comparisons on the Neu3DV dataset in Table 2. Relative to 4DGS, our method achieves up to a 125$\times$ compression ratio while preserving similar visual quality and rendering speed. It is observed that compared to the SOTA NeRF-based baseline MixVoxels (Wang et al., 2023), our method achieves a 20$\times$ storage reduction and a 16$\times$ inference speed improvement, maintaining comparable rendering quality. Furthermore, our approach exhibits higher rendering quality and smaller storage overhead compared to most Gaussian-based methods.

### v2

@@ 3.2 Memory-efficient 4D Gaussian Splatting
Overview. As illustrated in Fig. 3, we develop our memory-efficient 4D Gaussian framework to significantly reduce the number of per-Gaussian stored parameters and drive the model to reconstruct dynamic scene with fewer 4D Gaussians. During the rendering process, we utilize a set of optimized 4D Gaussians and initially transform each Gaussian based on specific time and view direction. This transformation procedure involves Gaussian color prediction and geometry deformation. By modifying the geometric structure of each Gaussian, we effectively broaden its action range. This expansion not only reduces the total number of Gaussians required but also increases the rendering participation rate of each Gaussian. Following the per-Gaussian transformation, we adhere to the established protocols of the original 4DGS [45] to carry out temporal slicing, projection, and differentiable rasterization, all critical for rendering high-quality frames.
Memory-efficient 4D Gaussian. 4DGS introduces 4D spherical harmonics $\boldsymbol{h}$ to model the temporal evolution of view-dependent color in dynamic scenes, which typically requires 144 of the total 161 parameters and contributes to the main storage overhead. While [20] have explored the use of a grid-based neural field to replace SH coefficients $\boldsymbol{h}$, we find that directly applying this method results in severe performance loss compared to 4DGS (see Table 3).
To overcome this issue, we propose a compact DC-AC color (DAC) representation. Specifically, we decouple the color attribute as a per-Gaussian DC color component $\boldsymbol{c}_{dc}\in\mathbb{R}^{3}$ and a temporal-viewpoint aware AC color predictor $\mathcal{F}_{\boldsymbol{\phi}}$. To predict the final color $\boldsymbol{c}_{t,v}$ of each Gaussian, we first compute the normalized view direction $\boldsymbol{d}_{v}=\frac{\boldsymbol{\mu}_{3D}-\boldsymbol{p}_{v}}{||\boldsymbol{\mu}_{3D}-\boldsymbol{p}_{v}||_{2}}$ for each Gaussian according to the camera center point $\boldsymbol{p}_{v}\in\mathbb{R}^{3}$ at the viewpoint $v$. Then, we concatenate the 3D position $\boldsymbol{\mu}_{3D}$, view direction $\boldsymbol{d}_{i}$, time $t$, and DC color $\boldsymbol{c}_{dc}$ and input them to a lightweight MLP network $\mathcal{F}_{\boldsymbol{\phi}}$:
$\boldsymbol{c}_{t,v}={\rm sigmoid}(\boldsymbol{c}_{dc}+\mathcal{F}_{\boldsymbol{\phi}}({\rm sg}(\boldsymbol{\mu}_{3D}),{\rm sg}(\boldsymbol{d}_{v}),t,\boldsymbol{c}_{dc})),$ (3)
where ${\rm sg}(\cdot)$ indicates a stop-gradient operation. This hybrid color composition method not only effectively preserves the individual information using DC component and supplements the missing viewpoint and time information using the AC predictor to maintain high rendering quality, but also reduces the storage overhead by up to 8×8$\times$8 × compared to the original 4DGS [45].
Table 1: Quantitative comparison with various competitive baselines on the Technicolor Dataset. “Storage" refers to the total model size for 50 frames.
[…]
DyNeRF [22] 31.80 - 0.0210 0.1400 0.02 30.00MB
HyperReel [1] 32.70 0.0470 - 0.1090 4.00 60.00MB
Deformable 3DGS [43] 30.95 0.0696 0.0353 0.1553 76.09 61.36MB
STG [23] 33.35 0.0404 0.0187 0.0846 141.73 51.35MB
E-D3DGS [2] 32.89 0.0494 0.0231 0.1114 79.14 56.07MB
4DGS [45] 32.07 0.0535 0.0263 0.1189 55.26 6107.07MB
Ours 33.57 0.0442 0.0204 0.1014 83.14 32.45MB
@@
Entropy-constrained Gaussian Deformation. For a specific time $t$, 4DGS [45] presupposes that the sliced 4D Gaussians exhibit linear movement while their rotation and scale remain constant. This strict assumption simplifies the movement representation and forces the model to combine multiple extra Gaussians to present any complex non-linear motions. Moreover, the sliced 4D Gaussian introduces the temporal decay opacity $\sigma_{t}$. From its definition, it is analyzed that a Gaussian gradually appears as time $t$ approaches its temporal position $\mu_{t}$, peaks in opacity at $t=\mu_{t}$, and gradually diminishes in density as $t$ moves away from $\mu_{t}$. As shown in Fig. 4 (a), this limited temporal operation range results in more than 90% of Gaussians being excluded at each time, causing the model to densify a large amount of Gaussians for rendering high-quality scene.
@@
To address these limitations, we advocate for improving flexibility in the motion representation and geometric structure of each 4D Gaussian. Specifically, we introduce a temporal-viewpoint aware deformation predictor to enlarge the action range of Gaussians. The 4D Gaussian center $\boldsymbol{\mu}_{4D}$, view direction $\boldsymbol{d}_{i}$, and time $t$ are mapped to a high-dimensional space using a regular frequency positional encoding function $\gamma$ [30], and then processed through a lightweight MLP network $\mathcal{F}_{\boldsymbol{\theta}}$ to predict the position deformation $\boldsymbol{m}_{\boldsymbol{\mu}_{4D}}^{t,v}\in\mathbb{R}^{4}$, scale deformation $\boldsymbol{m}_{\boldsymbol{s}_{4D}}^{t,v}\in\mathbb{R}^{4}$, and rotation deformations $\boldsymbol{m}_{\boldsymbol{q}_{l}}^{t,v}\in\mathbb{R}^{4},\boldsymbol{m}_{\boldsymbol{q}_{r}}^{t,v}\in\mathbb{R}^{4}$ as:
$(\boldsymbol{m}_{\boldsymbol{\mu}_{4D}}^{t,v},\boldsymbol{m}_{\boldsymbol{s}_{4D}}^{t,v},\boldsymbol{m}_{\boldsymbol{q}_{l}}^{t,v},\boldsymbol{m}_{\boldsymbol{q}_{r}}^{t,v})=$ (4)
$\mathcal{F}_{\boldsymbol{\theta}}(\gamma({\rm sg}(\boldsymbol{\mu}_{4D})),\gamma({\rm sg}(\boldsymbol{d}_{v})),\gamma(t)),$
@@
$\boldsymbol{\mu}_{4D}^{t,v}$ $=\boldsymbol{\mu}_{4D}\times\boldsymbol{m}_{\boldsymbol{\mu}_{4D}}^{t,v},\quad\boldsymbol{s}_{4D}^{t,v}=\boldsymbol{s}_{4D}\times\boldsymbol{m}_{\boldsymbol{s}_{4D}}^{t,v},$ (5)
$\quad\boldsymbol{q}_{l}^{t,v}$ $=\boldsymbol{q}_{l}\times\boldsymbol{m}_{\boldsymbol{q}_{l}}^{t,v},\quad\;\;\quad\boldsymbol{q}_{r}^{t,v}=\boldsymbol{q}_{r}\times\boldsymbol{m}_{\boldsymbol{q}_{r}}^{t,v}.$
@@ 3.3 Training and Compression Pipeline
Loss Function. Following the original 4DGS [45], we adopt the photometric loss $\mathcal{L}_{photo}$, consisting of $\mathcal{L}_{1}$ loss and structural similarity loss $\mathcal{L}_{ssim}$, to measure the distortion between the rendered image and ground truth image. By adding the loss for opacity regularization $\mathcal{L}_{opa}$, the overall loss $\mathcal{L}$ is defined as:
@@ 4.1 Experimental Setup
Datasets. We evaluate the effectiveness of our method using two real-world benchmarks that are representative of various challenges in dynamic scene rendering: (1) Technicolor Light Field Dataset [36]: This dataset consists of multi-view video data captured by a time-synchronized 4$\times$4 camera rig. Following HyperReel [1], we exclude the camera at the second row, second column and evaluate on five scenes (Birthday, Fabien, Painter, Theater, and Trains) at 2048$\times$1088 full resolution. (2) Neural 3D Video Dataset (Neu3DV) [22]: This dataset includes six indoor multi-view video scenes captured by 18 to 21 cameras, each at a resolution of 2704×2028 pixels. The scenes (Coffee Martini, Cook Spinach, Cut Roasted Beef, Flame Salmon, Flame Steak, Sear Steak) vary in duration and feature dynamic movements, some with multiple objects in motion. Consistent with existing practices [22, 45], evaluations are conducted at half resolution of 300-frame scenes.
Table 2: Quantitative comparisons with various competitive baselines on the Neural 3D Video Dataset. “Storage" refers to the total model size for 300 frames. 1: Only report the result on the Flame Salmon scene. 2: Exclude the Coffee Martini scene. 3: These methods train each model with a 50-frame video sequence to prevent memory overflow, requiring six models to complete the overall evaluation.
@@
Neural Volume1 [25] 22.80 - 0.0620 0.2950 - -
LLFF1 [29] 23.24 - 0.0200 0.2350 - -
DyNeRF1 [22] 29.58 - 0.0200 0.0830 0.015 28.00MB
HexPlane2,3 [3] 31.71 - 0.0140 0.0750 0.56 200.00MB
StreamRF [21] 28.26 - - - 10.90 5310.00MB
NeRFPlayer3 [38] 30.69 0.0340 - 0.1110 0.05 5130.00MB
HyperReel [1] 31.10 0.0360 - 0.0960 2.00 360.00MB
K-Planes [13] 31.63 - 0.0180 - 0.30 311.00MB
MixVoxels-L [40] 31.34 - 0.0170 0.0960 37.70 500.00MB
MixVoxels-X [40] 31.73 - 0.0150 0.0640 4.60 500.00MB
Dynamic 3DGS [28] 30.46 0.0350 0.0190 0.0990 460.00 2772.00MB
C-D3DGS [16] 30.46 - - 0.1500 118.00 338.00MB
Deformable 3DGS [43] 30.98 0.0331 0.0191 0.0594 29.62 32.64MB
E-D3DGS [2] 31.20 0.0259 0.0151 0.0304 69.70 40.20MB
STG3 [23] 32.04 0.0261 0.0145 0.0440 273.47 175.35MB
4DGS [45] 31.57 0.0290 0.0164 0.0573 96.69 3128.00MB
@@
Evaluation Metrics. To assess the quality of rendered videos, we utilize three popular image quality assessment metrics: Peak Signal-to-Noise Ratio (PSNR), Dissimilarity Structural Similarity Index Measure (DSSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [46]. PSNR quantifies the pixel color error between the rendered and original frames. DSSIM evaluates the perceived dissimilarity of the rendered image, while LPIPS measures the higher-level perceptual similarity using an AlexNet backbone [19]. Given the inconsistency in DSSIM implementation noted across different methods [13, 1], we follow [23] to distinguish DSSIM results into two categories: DSSIM1 and DSSIM2. DSSIM1 is calculated with a data range set to 1.0, based on the structural similarity function from the scikit-image library, whereas DSSIM uses a data range of 2.0. For rendering speed, we measure the performance in frames per second (FPS).
@@
Baselines. As we introduce MEGA, a novel method for compressing 4DGS [45], our primary comparison focuses on the baseline 4DGS method. Additionally, we benchmark MEGA against a range of NeRF-based baselines, including DyNeRF [22], HyperReel [1], Neural Volume [25], LLFF [29], HexPlane [3], StreamRF [21], NeRFPlayer [38], MixVoxels [40], and K-Planes [13]. Other recent competitive Gaussian-based methods are also considered in our comparisons, including Dynamic 3DGS [28], C-D3DGS [16], Deformable 3DGS [43], E-D3DGS [2], and STG [23]. The numerical results of Deformable 3DGS, E-D3DGS, STG, and 4DGS are produced by running their released codes on a single NVIDIA A800 GPU, while results for other baselines are from their original papers.
@@
Implementation Details. We train our MEGA model over 30k iterations and stop densification at the midpoint. We use the Adam optimizer with a batch size of one, adopting the hyperparameter settings from the original 4DGS [45] framework, including loss weight, learning rate, and threshold parameters. When rendering the view at time $t$, we filter out those Gaussians with $\sigma(t)\leq 0.05$. To ensure stable training of our deformation predictor, we introduce weight regularization and set it at $1e^{-6}$. The learning rate of the deformation predictor undergoes exponential decay, starting from $8e^{-4}$ and reducing to $1.6e^{-6}$. For the AC color predictor, we start with an initial learning rate of 0.01, incorporating a 100-step warm-up phase. Subsequently, its learning rate is decreased by a factor of three at the 5k, 15k, and 25k steps. Regarding the hyper-parameters in the loss function, we set $\lambda$ and $\kappa$ as 0.2 and 0.0005, respectively, to balance the contributions of different components.
Figure 5: Subjective comparison of various methods on Theater scene from the Technicolor Dataset.
@@
4.2 Experimental Results
Table 1 details a quantitative evaluation of our MEGA method on the Technicolor dataset. Notably, our method surpasses the main baseline 4DGS [45], with PSNR, DSSIM1, DSSIM2, and LPIPS improvements by 1.2dB, 0.01, 0.006, and 0.018, respectively. Meanwhile, it significantly reduces storage requirements, achieving a 190$\times$ compactness and improving rendering speed by 50%. When compared with the NeRF-based method HyperReel [1], MEGA achieves a substantial improvement in representation, with an increase of about 0.87dB in PSNR and a 20$\times$ faster rendering speed, while halving the storage overhead. Moreover, our MEGA records a 0.22dB gain in visual fidelity over the state-of-the-art Gaussian-based method STG [23], and reduces storage overhead by 40%. Fig. 5 offers qualitative comparisons for the Theater scene, demonstrating that our results contain more vivid details and provide artifact-less rendering.
Besides, we report the quantitative comparisons on the Neu3DV dataset in Table 2. Relative to 4DGS, our method achieves up to a 125$\times$ compression ratio while preserving similar visual quality and rendering speed. It is observed that compared to the SOTA NeRF-based baseline MixVoxels [40], our method achieves a 20$\times$ storage reduction and a 16$\times$ inference speed improvement, maintaining comparable rendering quality. Furthermore, our approach exhibits higher rendering quality and smaller storage overhead compared to most Gaussian-based methods.